A steam engine intake steam at 200°C and after doing work exhausts itdirectly in air at 100°C calculate the percentage of heat used for doingwork. Assume the engine to be an ideal engine?

Grade:11

1 Answers

Arun
6 years ago
Dear Vaibhav
 
T1 = 200°c = 473 K
T2 = 100°c = 373 K
n(efficiency) = W/Q1 = (T1 - T2)/T1
= 100/473 = 0.21
= 21%
 
Regards
Arun (askIITians forum expert)